# On Mirzakhani’s twist torus conjecture

### v2

@@ 5. Identifying orbit closures
Let $\widehat{\mathcal{L}}$ be an AIS in a stratum (^$\mathcal{H}(\widehat{}\underline{\kappa})$ of genus $\widehat{g}$ abelian differentials. Then if $\widehat{\mathcal{L}}$ has high rank, then either $\widehat{\mathcal{L}}$ is a component of (^$\mathcal{H}(\widehat{}\underline{\kappa})$ or the locus of all holonomy double covers of surfaces in a component of a stratum of quadratic differentials.
We would like to apply Theorem 5.2 to identify $\mathcal{L}\subset\mathcal{Q}^{1}\mathcal{M}_{g}(\underline{\kappa};\varepsilon)$, the $P$-orbit closure of a flat twist torus or horosphere slice. To do so, consider the locus ^(^$\widehat{\mathcal{L}}\subset\mathcal{H}(\widehat{}\underline{\kappa})$ of holonomy double covers, which is itself an AIS. Let $\underline{\kappa}_{\operatorname{odd}}$ denote those elements of $\underline{\kappa}$ that are odd.
If $\varepsilon=+1$ (hence #$\#\underline{\kappa}_{\operatorname{odd}}=0$), then $\widehat{q}\to q$ is the trivial double cover, $\widehat{\mathcal{Q}}\cong\mathcal{Q}$, and the rank of $\widehat{\mathcal{Q}}$ is $g$. Otherwise, we record the relevant numerics below.
@@
Let $\mathcal{Q}\subset\mathcal{Q}^{1}\mathcal{M}_{g}(\underline{\kappa};-1)$ be a component of a stratum and $\widehat{\mathcal{Q}}\subset\mathcal{H}(\widehat{}\underline{\kappa})$ the locus of holonomy double branched covers. Then ^$\widehat{}\underline{\kappa}$ is a partition of $22\widehat{g}-$ 2, where
@@
Proof.
This is [AW21, Lemma 4.2], but we give some details for completeness; see also [BW17, Theorem 26] for a related computation in the context of train tracks. The expression for $\widehat{g}$ can be computed using the Riemann–Hurwitz formula for the holonomy double cover $\widehat{q}\to q$ branched over the zeros of $q\in\mathcal{QM}(\underline{\kappa};-1)$. To compute rank, subtract the complex dimension of the kernel of the map to absolute cohomology from the complex dimension of $\mathcal{QM}(\underline{\kappa};-1)$ and divide by two. Since $\dim\mathcal{QM}(\underline{\kappa};-1)=2g-2+\#\underline{\kappa}$ and the kernel of the map to absolute cohomology is generated by certain relative cycles associated to the even order zeros of any $q\in\mathcal{Q}$, the formula follows. ∎
Corollary 5.4.
Suppose that $\mathcal{L}\subset\mathcal{Q}^{1}\mathcal{M}_{g}(\underline{\kappa};-1)$ is an AIS and that
$\text{rank}(\mathbb{R}\widehat{\mathcal{L}})\geq g+\frac{1}{4}\#\underline{\kappa}_{\operatorname{odd}}+\frac{1}{2}.$
Then $\mathcal{L}$ is a component of $\mathcal{Q}^{1}\mathcal{M}_{g}(\underline{\kappa};-1)$.
Proof.
Clearly $\widehat{\mathcal{L}}$ is contained inside the locus of all holonomy double covers of differentials in $\mathcal{Q}^{1}\mathcal{M}_{g}(\underline{\kappa};-1)$. The rank hypothesis together with Lemma 5.3 implies that
$\text{rank}(\widehat{\mathcal{L}})\geq g+\frac{1}{4}\#\underline{\kappa}_{\operatorname{odd}}+\frac{1}{2}=\widehat{g}/2+1,$
hence $\mathbb{R}\widehat{\mathcal{L}}$ has high rank and so we can apply Theorem 5.2. ∎
Remark 5.5.
We note that, by the second computation in Lemma 5.3, the hypotheses of the Corollary can only be satisfied when there are at least 6 odd order zeros.
5.1. Cylinder deformations and rank
To apply Theorem 5.2, we must bound the rank of $\mathcal{L}$ containing a twist torus or a horosphere slice from below. For this, we use cylinder deformations, which were introduced in [Wri15] and are recalled below.
For a given quadratic differential $q$, let $\mathcal{C}$ be the set of maximal, singularity free horizontal cylinders in $q$. Each element $C_{i}\in\mathcal{C}$ is an isometrically embedded open cylinder with height $h_{i}>0$ and circumference $\ell_{i}>0$, and $C_{i}$ is foliated by closed nonsingular leaves of the imaginary foliation of $q$ which are all homotopic to its core curve $\gamma_{i}$. Recall that
$u_{s}=\begin{pmatrix}1&s\\ 0&1\end{pmatrix}\in\operatorname{\mathsf{PSL}}_{2}\mathbb{R},$
and define $u_{s}^{C_{i}}q$ to be the holomorphic quadratic differential obtained by applying the horocycle flow $u_{s}$ to $C_{i}$ but not the rest of $q$.
For the rest of this section, $\Theta_{\gamma}\subset\mathcal{L}$ denotes either a (flat) twist torus $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ or a horosphere (slice) $\mathcal{Q}H_{\gamma}(\mathbf{h},\Gamma)$ as constructed in §4. Then each $q\in\Theta_{\gamma}$ is completely horizontally periodic, the core curves of the horizontal cylinders $\mathcal{C}$ constitute $\gamma=\cup_{i=1}^{n}\gamma_{i}$, and the height of the cylinder $C_{i}$ corresponding to $\gamma_{i}$ is $h_{i}$. From the definitions, we see that
$u_{s_{1}}^{C_{1}}\cdot...\cdot u_{s_{n}}^{C_{n}}q\in\Theta_{\gamma}$
for all $(s_{1},...,s_{n})\in\mathbb{R}^{n}$. Denote by $\eta_{i}$ the tangent vector to the path $s\mapsto u_{s}^{C_{i}}q$ at $s=0$, so that
(6) $\langle\eta_{1},...,\eta_{n}\rangle\leq T_{q}\Theta_{\gamma}\leq T_{q}\mathcal{L}\cong T_{\widehat{q}}\widehat{\mathcal{L}}.$
The linear holonomy along the core curve of a horizontal cylinder is trivial, so we know the preimage of $\gamma_{i}$ in the holonomy cover $\widehat{q}$ has two components. Pick a lift $\overline{\gamma}_{i}\subset\widehat{q}$ so that the full preimage is $\overline{\gamma}_{i}\sqcup\iota(\overline{\gamma}_{i})$, where $\iota$ denotes the covering involution. Choose an orientation on each $\gamma_{i}\subset S$ and orient $\overline{\gamma}_{i}$ so that it maps with degree $1$ to $\gamma_{i}$. Push that orientation forward by $\iota$ to induce an orientation on $\iota(\overline{\gamma}_{i})$ and define $\widehat{\gamma}_{i}=\overline{\gamma}_{i}-\iota_{*}\overline{\gamma}_{i}\in H_{1}(\widehat{q}\setminus Z(\widehat{q});\mathbb{Z})$. It satisfies $\iota_{*}\widehat{\gamma}_{i}=-\widehat{\gamma}_{i}$. See Figure 6.
The following Lemma is the content of [Wri15, §2] (after lifting the situation from $q$ to $\widehat{q}$).
Lemma 5.6.
Up to multiplication by ±$\pm 1$, the deformation $\eta_{i}\in H^{1}(\widehat{q},Z(\widehat{q});\mathbb{C})^{-}\cong T_{\widehat{q}}\mathbb{R}\widehat{\mathcal{Q}}$ is Poincaré–Lefschetz dual to the cylinder $h_{i}\widehat{\gamma}_{i}\in H_{1}(\widehat{q}\setminus Z(\widehat{q});\mathbb{C})^{-}$.
The inclusion induces a natural map $i_{*}:H_{1}(\widehat{q}\setminus Z(\widehat{q});\mathbb{Z})\to H_{1}(\widehat{q};\mathbb{Z})$ and Poincaré–Lefschetz duality makes the following square commute (with any coefficients):
@@ 5.1. Cylinder deformations and rank
Proposition 5.7.
The rank of any AIS $\mathcal{L}$ containing $\Theta_{\gamma}$ is bounded below by the real dimension of
$i_{*}\langle\widehat{\gamma}_{1},...,\widehat{\gamma}_{n}\rangle_{\mathbb{R}}\leq H_{1}(\widehat{q};\mathbb{R})^{-}.$
Proof.
Lemma 5.6 together with (6) give that
(7) $p(\langle\eta_{1},...,\eta_{n}\rangle)=i_{*}\langle\widehat{\gamma}_{1},...,\widehat{\gamma}_{n}\rangle_{\mathbb{R}}\otimes\mathbb{C}.$
By [AEM17], $p(T_{\widehat{q}}\widehat{\mathcal{L}})$ is symplectic, and since $\gamma_{i}\cap\gamma_{j}=\emptyset$ for $i\not=j$, the subspace (7) is isotropic. Thus the complex dimension of $p(T_{\widehat{q}}\widehat{\mathcal{L}})$ is bounded below by twice the complex dimension of (7). ∎
5.2. Lifting homologies
Proposition 5.7 reduces our problem to a topological question about (the lift of) $\gamma$ in the surface $\widehat{S}$ homeomorphic to $\widehat{q}$. The topology of $\widehat{S}$ depends not just on $\gamma$, but also on the geometry of the complement of $\gamma$ in the corresponding hyperbolic/flat structures. The relevant data is encoded in the geometric arc system $\underline{\alpha}$ dual to the ribbon graph $\mathsf{Y}$ or $\Gamma$.
Definition 5.8.
Let $\Sigma$ be a compact surface with non-empty boundary and let $\underline{\alpha}$ be a filling arc system on $\Sigma$. Say that $\underline{\alpha}$ is co-orientable on $\Sigma$ if there is an orientation of $\partial\Sigma$ and a co-orientation of $\underline{\alpha}$ continuously extending the orientation on $\partial\Sigma$.
Given a multicurve $\gamma\subset S$ and a filling arc system $\underline{\alpha}$ on $S\setminus\gamma$, the pair $(\gamma,\underline{\alpha})$ is jointly orientable if $\underline{\alpha}$ can be co-oriented on all complementary subsurfaces inducing a consistent orientation on $\gamma$.
Clearly, if $(\gamma,\underline{\alpha})$ is jointly orientable, then the restriction of $\underline{\alpha}$ to each component of $S\setminus\gamma$ is co-orientable. On the other hand, $(\gamma,\underline{\alpha})$ need not be jointly orientable even if the restriction of $\underline{\alpha}$ to each component of $S\setminus\gamma$ is co-orientable. See Figure 4.
Figure 4. A co-orientable $\underline{\alpha}$ on $S\setminus\gamma$ where $(\gamma,\underline{\alpha})$ is not jointly orientable.
Co-orientability and joint orientability are related to the topology of $\widehat{S}\to S$ as follows.
Lemma 5.9.
Let $q\in\Theta_{\gamma}$ and consider the corresponding holonomy double branched cover $\widehat{S}\to S$. For any component $\Sigma\subset S\setminus\gamma$, the preimage $\widehat{\Sigma}$ is disconnected if and only if $\underline{\alpha}|_{\Sigma}$ is co-orientable, and $\widehat{S}$ is disconnected if and only if $(\gamma,\underline{\alpha})$ is jointly orientable.
Proof.
We begin by observing that we can fix nice representatives for the components of $S\setminus\gamma$ and for $\underline{\alpha}$ as follows. By construction, each curve of $\gamma$ can be realized as a core curve of a horizontal cylinder of $q$. Cutting along these core curves results in a flat cone surface with boundary homeomorphic to $S\setminus\gamma$. Each component $\Sigma$ obtained this way deformation retracts onto the graph of horizontal saddle connections contained in $\Sigma$ by collapsing the segments of its vertical foliation, and we can pick a representative of $\underline{\alpha}$ by choosing one segment in each isotopy class. Compare Figure 5.
From this description, it follows that $\underline{\alpha}$ is co-orientable on $\Sigma$ if and only if the restriction of the horizontal foliation $\operatorname{Im}(q)$ to $\Sigma$ is orientable (see also the discussion in Section 7.2 of [CF24b]), which happens if and only if the preimage $\widehat{\Sigma}$ is disconnected. Similarly, a joint orientation on $(\gamma,\underline{\alpha})$ is equivalent to a global choice of orientation of $\operatorname{Im}(q)$ on $S$, which implies the holonomy double cover is disconnected. ∎
Figure 5. A co-orientation on $\underline{\alpha}|_{\Sigma}$ and the corresponding orientation of the horizontal foliation of $q$ restricted to $\Sigma$.
The following Proposition is proved by a technical homology computation in $\widehat{S}$. The basic idea is that each subsurface of $S\setminus\gamma$ defines a linear relation between its boundary curves in homology on $S$, and these relations lift to relations between the curves of $\widehat{\gamma}$ in $\widehat{S}$ exactly when that subsurface is disconnected in $\widehat{S}$. The main difficulty in the proof comes from ensuring that there are no other unaccounted-for relations.
Proposition 5.10.
Any invariant subvariety $\mathcal{L}\subset\mathcal{Q}^{1}\mathcal{M}_{g}$ containing $\Theta_{\gamma}$ with defining arc system $\underline{\alpha}$ (dual to $\mathsf{Y}$ or $\Gamma$) has rank at least
#$\#\gamma-N_{co}+\delta_{jo},$
where $N_{co}$ is the number of components $\Sigma$ of $S\setminus\gamma$ for which $\underline{\alpha}|_{\Sigma}$ is co-orientable, and $\delta_{jo}=1$ if $(\gamma,\underline{\alpha})$ is jointly orientable and 00 otherwise.
Proof.
We will show that #$\#\gamma-N_{co}+\delta_{jo}$ is precisely the dimension of the space
$V=i_{*}\langle\widehat{\gamma}_{1},...,\widehat{\gamma}_{n}\rangle\leq H_{1}(\widehat{S};\mathbb{R})^{-},$
so that Proposition 5.7 establishes the relationship with rank. For brevity, and to avoid confusion with the action $\iota_{*}$ of the covering involution on the homology of $\widehat{S}$, we drop the $i_{*}$ and write $\widehat{\gamma}_{i}\in H_{1}(\widehat{S};\mathbb{R})^{-}$.
@@ 5.2. Lifting homologies
Enumerate the subsurfaces $S\setminus\gamma=\Sigma_{1}\cup...\cup\Sigma_{k}$, and let $\widehat{\Sigma}_{i}$ be the full preimage of $\Sigma_{i}$ under $\widehat{S}\to S$. Consider annular neighborhoods $A_{i}$ of $\gamma_{i}\subset S$ and let $\widehat{A}_{i}$ be the full preimages. Recall that the anti-invariant part of $H_{1}(\widehat{A}_{i})$ is generated by $\widehat{\gamma}_{i}=\overline{\gamma}_{i}-\iota_{*}\overline{\gamma}_{i}$.
Figure 6. A non-trivial cover $\widehat{\Sigma}\to\Sigma$ corresponding to an arc system $\underline{\alpha}|_{\Sigma}$ that is not co-orientable. We have $\widehat{\gamma}_{i}=\overline{\gamma}_{i}-\iota_{*}\overline{\gamma}_{i}$ and the push-off to the right is $\widehat{\gamma}_{i}^{+}=\overline{\gamma}_{i}^{+}-\iota_{*}\overline{\gamma}_{i}^{+}$.
By Lemma 5.9, a subsurface $\widehat{\Sigma}_{j}$ is disconnected if and only if $\underline{\alpha}|_{\Sigma_{j}}$ is co-orientable. Our claim is then that the relation among the boundary curves of $\Sigma_{j}$ in homology yields a relation among the corresponding homology classes in $\widehat{\Sigma}_{i}$ when $\underline{\alpha}|_{\Sigma_{j}}$ is co-orientable, and otherwise the relation does not lift. We also claim that these are the only relations satisfied by the $\{\widehat{\gamma}_{i}\}$, and that there is even a redundancy among them in the case that $(\gamma,\underline{\alpha})$ is jointly orientable.
@@
To make this precise, consider the Mayer–Vietoris long exact sequence (with coefficients in $\mathbb{R}$, say)
(8) $0\to H_{2}(\widehat{S})\to\oplus_{i,j}H_{1}(\widehat{A}_{i}\cap\widehat{\Sigma}_{j})\to\left(\oplus_{i=1}^{n}H_{1}(\widehat{A}_{i})\right)\bigoplus\left(\oplus_{j=1}^{k}H_{1}(\widehat{\Sigma}_{j})\right)\to H_{1}(\widehat{S})\to...$
@@
Assume first that $\widehat{S}$ is connected, i.e., that $(\gamma,\underline{\alpha})$ is not jointly orientable (Lemma 5.9). Since $\iota$ is orientation-preserving and $\widehat{S}$ is closed, we have $H_{2}(\widehat{S})^{-}=0$. From the definitions, $H_{1}(\widehat{A}_{i})^{-}=\langle\widehat{\gamma}_{i}\rangle$. Let $\gamma_{i}^{+}$ and $\gamma_{i}^{-}$ be curves parallel to $\gamma_{i}$ obtained by pushing off to the right and left, respectively. For any fixed $i$, we have that $\bigcup_{j}\widehat{A}_{i}\cap\widehat{\Sigma}_{j}$ is a union of four cylinders, the cores of which correspond to $\overline{\gamma}_{i}^{\pm}$ and $\iota_{*}\overline{\gamma}_{i}^{\pm}$. Thus $\oplus_{i,j}H_{1}(\widehat{A}_{i}\cap\widehat{\Sigma}_{j})^{-}$ is freely generated by $\{\widehat{\gamma}_{i}^{\pm}:i=1,...,n\}$.
Recall that $\widehat{\Sigma}$ is the total preimage of $\Sigma$ under the two-fold branch covering projection, which is then invariant under the covering involution. Hence the anti-invariance distributes over the sum in (8), yielding
@@
where $F:\widehat{\gamma}_{i}^{\pm}\mapsto\widehat{\gamma}_{i}-\widehat{\gamma}_{i}^{\pm}$ and $G$ sums the components together. Note $\dim\operatorname{im}F=2n$ since (9) is exact.
@@
is a linearly independent set if and only if $\widehat{\Sigma}_{j}$ is connected and otherwise spans a subspace of dimension $m_{j}-1$. By definition, there are $N_{co}$ such disconnected $\widehat{\Sigma}_{j}$, so that $\dim G^{-1}(V)=3n-N_{co}$.
Since $\widehat{S}$ is connected by assumption, i.e., $(\gamma,\underline{\alpha})$ is not jointly orientable and $\delta_{jo}=0$, we have that $V\cong G^{-1}(V)/\operatorname{im}F$ has dimension
$\dim G^{-1}(V)-\dim\operatorname{im}F=(3n-N_{co})-2n=n-N_{co}+0=\#\gamma-N_{co}+\delta_{jo}.$
In the case that $\widehat{S}$ is disconnected, i.e., $\delta_{jo}=1$, the computation is identical except that $H_{2}(\widehat{S})^{-}\cong H_{2}(S)$, and so $\operatorname{im}F$ has dimension $12n-$ 1. Thus $V$ has dimension $3n-N_{co}-(2n-1)=\#\gamma-N_{co}+\delta_{jo}$, completing the proof of the Proposition. ∎
5.3. Orbit closures of twist tori
We now apply the computation of the previous subsection. Let $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ be a flat twist torus constructed from a multicurve $\gamma$ (not necessarily a pants decomposition). Let $\underline{\alpha}$ be the arc system dual to $\mathsf{Y}$, and let $\underline{\kappa}$ be the partition of $44g-$ 4 determined by the valences of the vertices of $\mathsf{Y}$. Let $\varepsilon=1$ if $(\gamma,\underline{\alpha})$ is jointly orientable and $\varepsilon=-1$, otherwise.

### v3

@@ 6.2. Jointly orientable constructions
Any high rank AIS in a stratum of squares of abelian differentials is either a component of the stratum or the locus of all holonomy double covers of surfaces in a component of a stratum of quadratic differentials.
Given any twist torus $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ contained in such a stratum of squares of abelian differentials, the dimension of the projection of the tangent space of the twist torus to absolute cohomology (as in Proposition 5.4) is just the dimension of the span of the curves of $\gamma$ in homology. Thus, so long as $\gamma$ has at least $g/2+1$ homologically independent curves, the $P$ orbit closure of $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{Y},\mathbf{h})$ is the unit area locus in either the ambient stratum or a locus of double covers.
Combining this Theorem together with the argument from Section 5.2, mutatis mutandis, immediately yields a result for jointly orientable gluings of regular plumbing fixtures.
Theorem 6.2.
Let $\gamma$ be a multicurve on a surface $S$ such that $S\setminus\gamma$ is a union of $p_{1},\ldots,p_{k}$-holed spheres, where all $p_{i}$ are even. Fix $L>0$ and consider the twist tori in $\mathcal{M}_{g}$ obtained by gluing together $p_{i}$-holed regular plumbing fixtures with all boundary lengths $L$ along $\gamma$. Suppose furthermore that $(\gamma,\mathsf{X})$ is jointly orientable, where $\mathsf{X}=\mathsf{X}(p_{1},\ldots,p_{k})$ is the ribbon graph spine for the regular plumbing fixtures. Fix any choice of heights $\mathbf{h}$ such that $\sum h_{i}=1/L$. There is a set of times $Z$ of zero density such that
[…]
  1. (1)
    If $S$ admits an involution fixing each of the curves of $\gamma$ setwise but reversing their orientations (hence $k\leq 2$), then $\mathcal{L}$ is the hyperelliptic locus in the component of the stratum
    $\mathcal{QM}_{g}\left((p_{1}-2)^{2},\ldots,(p_{k}-2)^{2};+1\right).$
  2. (2)
    Otherwise, $\mathcal{L}$ is the entire component of the ambient stratum containing $\mathcal{Q}\mathbb{T}_{\gamma}(\mathsf{X},\mathbf{h})$.
To identify the $P$-orbit closure of the twist tori in this theorem, it suffices to use the full rank result of Mirzakhani–Wright [MW18], rather than Theorem 6.1.
Proof.
We note that since $\gamma$ cuts $S$ into surfaces without genus, it spans a $g$-dimensional subspace of homology. In particular, the orbit closure of the twist torus is high (in fact, full) rank. The only thing that remains to be justified is why no other loci of double covers of quadratic differentials can appear. For this to be the case, $\gamma$ must be setwise invariant under some orientation-preserving involution $\iota$ of $S$. Unless $\iota$ is hyperelliptic, there are two curves of $\gamma$ which must be interchanged. Twisting those curves independently is not $\iota$-invariant, hence the tangent space to the twist torus is not $\iota$-invariant. ∎
This proof is clearly much more general: one can pick any metric ribbon graph $\mathsf{X}$ such that $(\gamma,\mathsf{X})$ is jointly orientable and get the same result. See Figure 6 for examples that land in the minimal strata.
Remark 6.3.
Unlike Theorem 1.2, the gluing pattern of $\gamma$ can determine the limiting measure, as some strata $\mathcal{Q}\mathcal{M}_{g}(\underline{\kappa};+1)$ are disconnected. However, one can read off the parity of the induced spin structure from the data of $\gamma$ and $\mathsf{X}$ by computing winding numbers.
Figure 6. Twist tori corresponding to the minimal stratum of squares of abelian differentials. The drawn arc systems are dual to ribbon graphs each with a unique vertex of valence 4g24g-2. The leftmost torus lands in the hyperelliptic locus, while the other two land in the odd and even spin components (which diagram corresponds to which spin depends on the parity of $g$ mod 44, see [CS21, §6.3]).
Using more specific results from Teichmüller dynamics, we can identify the limiting distributions of more twist tori. We mention in particular the following result of Winsor [Win24], which itself admits further generalizations.
Theorem 6.4 (Theorem 1.3 of [Win24]).
Suppose $\mathcal{L}$ is an AIS in a component $\mathcal{H}$ of a stratum of squares of abelian differentials. Suppose further there is some $\omega\in\mathcal{L}$ with a set of three horizontal cylinders whose core curves bound a pair of pants, and that cylinder deformations on each of these cylinders remains in $\mathcal{L}$. Then $\mathcal{L}=\mathcal{H}$.
Consequently, if $\gamma$ is any multicurve such that some component of $S\setminus\Gamma$ is a pair of pants, and $\mathsf{Y}\in\mathcal{T}(S\setminus\gamma)$ is any metric ribbon graph such that $(\gamma,\mathsf{Y})$ is jointly orientable, then the twist tori $\mathcal{P}\mathbb{T}_{\gamma}(e^{t}\mathsf{Y},e^{-t}\mathbf{h})$ equidistribute outside a set of zero density to the pullback of Masur–Smillie–Veech measure coming from their ambient stratum.
6.3. Lifting homologies
When $(\gamma,\mathsf{Y})$ is not jointly orientable we have fewer tools at our disposal, as the classification of orbit closures in quadratic strata is not as far along as the abelian case (for example, there is no known analogue of Theorem 6.4). However, we can still apply Theorem 6.1 in many situations.
In order to do this, let us record some numerical invariants of quadratic doubles. Let $\mathcal{Q}$ be a component of a stratum $\mathcal{QM}_{g}(\underline{\kappa};-1)$ of genus $g$ holomorphic quadratic differentials, where $\underline{\kappa}$ is a partition of 4g44g-4. Let $\underline{\kappa}_{\operatorname{odd}}$ denote those elements of $\underline{\kappa}$ that are odd and likewise define $\underline{\kappa}_{\textrm{even}}$.
@@ 6.3. Lifting homologies
Let $\mathcal{Q}\subset\mathcal{Q}\mathcal{M}_{g}(\underline{\kappa};-1)$ be a component of a stratum and let $\widehat{\mathcal{Q}}\subset\mathcal{Q}\mathcal{M}_{\widehat{g}}(\widehat{}\underline{\kappa};+1)$ be the AIS consisting of holonomy double covers of surfaces in $\mathcal{Q}$. Then
$\widehat{g}=2g+\frac{1}{2}\#\underline{\kappa}_{\operatorname{odd}}-1,\hskip 17.22217pt\textrm{rank}(\widehat{\mathcal{Q}})=g+\frac{1}{2}\#\underline{\kappa}_{\operatorname{odd}}-1,\hskip 17.22217pt\text{and}\hskip 17.22217pt\textrm{rel}(\widehat{\mathcal{Q}})=\#\underline{\kappa}_{\textrm{even}}.$
Proof.
This is [AW21, Lemma 4.2], but we give some details for completeness; see also [BW17, Theorem 26] for a related computation in the context of train tracks. The expression for $\widehat{g}$ can be computed using the Riemann–Hurwitz formula for the holonomy double cover $\hat{q}\to q$ branched over the zeros of $q\in\mathcal{QM}(\underline{\kappa};-1)$. To compute rank, subtract the complex dimension of the kernel of the map to absolute cohomology from the complex dimension of $\mathcal{QM}(\underline{\kappa};-1)$ and divide by two. Since $\dim\mathcal{QM}(\underline{\kappa};-1)=2g-2+\#\underline{\kappa}$ and the kernel of the map to absolute cohomology is generated by certain relative cycles associated to the even order zeros of any $q\in\mathcal{Q}$, the formulas follow. ∎
Remark 6.6.
We note that the locus of double covers $\widehat{\mathcal{Q}}$ is only high rank in $\mathcal{Q}\mathcal{M}_{\widehat{g}}(\widehat{}\underline{\kappa};+1)$ once $\underline{\kappa}_{\operatorname{odd}}\geq 6.$
Proposition 5.4 gives a lower bound for the rank of an AIS $\mathcal{L}$ in terms of cylinder deformations. When $\mathcal{L}$ contains the entire twist torus $\mathcal{Q}\mathbb{T}(\mathsf{Y},\mathbf{h})$, this dimension can be computed explicitly.
Proposition 6.7.
Let $\mathcal{Q}$ be a component of $\mathcal{QM}_{g}(\underline{\kappa};-1)$. Let $q\in\mathcal{Q}$ be horizontally periodic with horizontal multicurve $\gamma$ and let $\Gamma$ denote the ribbon graph of horizontal separatrices on $q$. Then
$\dim_{\mathbb{R}}p(\mathrm{Twist}(\hat{q},\widehat{\mathcal{Q}}))=\#\gamma-N_{o}(\Gamma),$
where $N_{o}(\Gamma)$ is the number of orientable components of $\Gamma$.
Proof.
The basic idea is that cylinder deformations are dual to their core curves, and each subsurface of $S\setminus\gamma$ defines a linear relation between its boundary curves in the homology of $S$. These relations lift to relations between the curves of $\widehat{\gamma}$ in the homology double cover $\widehat{S}$ exactly when that subsurface is disconnected in $\widehat{S}$. The main technical difficulty is just to ensure that there are no other unaccounted-for relations.
Let us begin by setting some notation. Enumerate the curves of $\gamma$ as $\gamma_{1},\ldots,\gamma_{n}$. The linear holonomy along the core curve of a horizontal cylinder is trivial, so we know the preimage of each $\gamma_{i}$ in the holonomy cover $\hat{q}$ has two components. Pick a lift $\overline{\gamma}_{i}\subset\hat{q}$, so the full preimage is $\overline{\gamma}_{i}\sqcup\iota(\overline{\gamma}_{i})$, where $\iota$ denotes the covering involution. Choose an orientation on each $\gamma_{i}\subset S$ and orient $\overline{\gamma}_{i}$ so that it maps with degree $1$ to $\gamma_{i}$. Push that orientation forward by $\iota$ to induce an orientation on $\iota(\overline{\gamma}_{i})$ and define
$).\widehat{\gamma}_{i}=\overline{\gamma}_{i}-\iota_{*}\overline{\gamma}_{i}\in H_{1}(\hat{q}\setminus Z(\hat{q}$$\mathbb{Z})$.
Up to scale, these are the Poincaré-Lefschetz duals of the lifts of cylinder deformations of $q$ to $\hat{q}$, and satisfy $\iota_{*}\widehat{\gamma}_{i}=-\widehat{\gamma}_{i}$. See Figure 7.
@@
${H^{1}(\hat{q},Z(\hat{q}))^{-}}$${H_{1}(\hat{q}\setminus Z(\hat{q}))^{-}}$${H^{1}(\hat{q})^{-}}$${H_{1}(\hat{q})^{-}}$$\scriptstyle{p}$$\scriptstyle{i_{*}}$
Thus, in order to prove the Proposition it is equivalent to show
$\dim_{\mathbb{R}}i_{*}\langle\widehat{\gamma}_{1},...,\widehat{\gamma}_{n}\rangle=n-N_{co}.$
@@
Enumerate the subsurfaces $S\setminus\gamma=\Sigma_{1}\cup...\cup\Sigma_{k}$, and let $\widehat{\Sigma}_{j}$ be the full preimage of $\Sigma_{j}$ under $\widehat{S}\to S$. Let $\Gamma_{j}$ denote the component of $\Gamma$ corresponding to $\Sigma_{j}$, and observe that a subsurface $\widehat{\Sigma}_{j}$ is disconnected if and only if $\Gamma_{j}$ is orientable (as in the proof of Lemma 4.2). Consider annular neighborhoods $A_{i}$ of $\gamma_{i}\subset S$ and let $\widehat{A}_{i}$ be their full preimages.
@@
Now consider the Mayer–Vietoris long exact sequence (with coefficients in $\mathbb{R}$, say)
[…]
@@
Since $\iota$ is orientation-preserving and $\widehat{S}$ is closed, we have $H_{2}(\widehat{S})^{-}=0$. By definition, $H_{1}(\widehat{A}_{i})^{-}=\langle\widehat{\gamma}_{i}\rangle$. Let $\gamma_{i}^{+}$ and $\gamma_{i}^{-}$ be curves parallel to $\gamma_{i}$ obtained by pushing off to the right and left, respectively; see Figure 7. For any fixed $i$, we have that $\bigcup_{j}\widehat{A}_{i}\cap\widehat{\Sigma}_{j}$ is a union of four cylinders, the cores of which correspond to $\overline{\gamma}_{i}^{\pm}$ and $\iota_{*}\overline{\gamma}_{i}^{\pm}$. Thus $\oplus_{i,j}H_{1}(\widehat{A}_{i}\cap\widehat{\Sigma}_{j})^{-}$ is freely generated by $\{\widehat{\gamma}_{i}^{\pm}:i=1,...,n\}$.
Figure 7. A non-trivial cover $\widehat{\Sigma}\to\Sigma$ corresponding to a non-orientable ribbon graph. We have $\widehat{\gamma}_{i}=\overline{\gamma}_{i}-\iota_{*}\overline{\gamma}_{i}$ and the push-off to the right is $\widehat{\gamma}_{i}^{+}=\overline{\gamma}_{i}^{+}-\iota_{*}\overline{\gamma}_{i}^{+}$.
Since each $\widehat{\Sigma}_{j}$ is the total preimage of $\Sigma_{j}$, it is invariant under the covering involution $\iota_{*}$. Hence the anti-invariance distributes over the sum in (6), yielding
@@
where $F:\widehat{\gamma}_{i}^{\pm}\mapsto(\widehat{\gamma}_{i},-\widehat{\gamma}_{i}^{\pm})$ and $G$ sums the components together. Note $\dim(\operatorname{im}F)=2n$ since (7) is exact.
@@
is a linearly independent set if and only if $\widehat{\Sigma}_{j}$ is connected and otherwise spans a subspace of dimension $m_{j}-1$. By definition, there are $N_{o}(\Gamma)$ disconnected $\widehat{\Sigma}_{j}$, so that $\dim G^{-1}(V)=3n-N_{o}(\Gamma)$. Putting this all together, $V\cong G^{-1}(V)/\operatorname{im}F$ has dimension
$\dim G^{-1}(V)-\dim(\operatorname{im}F)=(3n-N_{o}(\Gamma))-2n=n-N_{o}(\Gamma),$
completing the proof of the Proposition. ∎
Combining Lemma 6.5 and Proposition 6.7 gives a criterion for when an AIS $\mathcal{L}$ containing a twist torus must have high rank. Pugging this into Theorem 6.1 yields the following: